OpenStudy (rayford):

How many times will the graph of y = –x intersect the graph of y = x2 – 3 ?

Parth (parthkohli):

Actually, you can equate \(-x\) and \(x^2 - 3\)

OpenStudy (lgbasallote):

equate them to one another \[x^2 - 3 = x\] now solve for x \[x^2 + x - 3 = 0\] now use the discriminant to find out how many positive roots are there \[b^2 - 4ac = 1^2 - 4(1)(-3) \implies 1 + 12 \implies 13\] it's positive therefore there are two roots. therefore it crosses twice

OpenStudy (lgbasallote):

got it?

Parth (parthkohli):

Since the solution sets will be the same, the point will be the same.. and so we'd have an intersection of those functions.

Parth (parthkohli):

Why \(x^2 - 3 = x\)? Why not \(x^2 - 3 = -x\)?

OpenStudy (lgbasallote):

uhh typo @_@

OpenStudy (lgbasallote):

\[x^2 - 3 = -x\] solve for x \[x^2 + x - 3 = 0\] now use discriminant to find how many roots are there \[b^2 - 4ac \implies (1)^2 - 4(-3)(1) \implies 1 + 12 \implies 13\] positive so there are 2 roots therefore crosses twice

OpenStudy (rayford):

thanks alot! u shold be a teacher man!

OpenStudy (lgbasallote):

i am haha

OpenStudy (rayford):

oh....well

OpenStudy (rayford):

you are good!